A cowboy fires a silver bullet of mass 3 g with a muzzle speed of 162 m/s into the pine wall of a saloon. What is the temperature change of the bullet? Assume that all the internal energy generated by the impact remains with the bullet. The specific heat of silver is 234 J/kg ·◦ C. Answer in units of ◦C.
Answer #1

kinetic energy of bullet = 0.5*m*v2 = 0.5*0.003*1622 = 39.366 joules

let temperature change be T

0.003*234*T=39.366

T=56.1 oC

Therefore, temperature increases by 56.1 oC